# Source Code Vulnerability Detection: Combining Code Language Models and Code Property Graphs

### v1

@@ Abstract
Currently, deep learning successfully applies to code vulnerability detection by learning from code sequences or property graphs. However, sequence-based methods often overlook essential code attributes such as syntax, control flow, and data dependencies, whereas graph-based approaches might underestimate the semantics of code and face challenges in capturing long-distance contextual information.
To address this gap, we propose Vul-LMGNN, a unified model that combines pre-trained code language models with code property graphs for code vulnerability detection. Vul-LMGNN constructs a code property graph that integrates various code attributes (including syntax, flow control, and data dependencies) into a unified graph structure, thereafter leveraging pre-trained code model to extract local semantic features as node embeddings in the code property graph. Furthermore, to effectively retain dependency information among various attributes, we introduce a gated code Graph Neural Network (GNN). By jointly training the code language model and the gated code GNN modules in Vul-LMGNN, our proposed method efficiently leverages the strengths of both mechanisms. Finally, we utilize a pre-trained CodeBERT as an auxiliary classifier, with the final detection results derived by learning the linear interpolation of Vul-LMGNN and CodeBERT. The proposed method, evaluated across four real-world vulnerability datasets, demonstrated superior performance compared to six state-of-the-art approaches. Our source code could be accessed via the link: https://github.com/Vul-LMGNN/vul-LMGGNN.
I Introduction
With the rapid expansion of the open-source community, software vulnerability detection technology has become a significant concern in the software industry and cybersecurity domain. Vulnerabilities pose a threat to the integrity and availability of software and computer systems, potentially leading to privilege escalation, leakage of sensitive data, denial of service, and various other attacks, resulting in substantial economic and societal losses [1]. In practice, developers and security engineers primarily rely on code analysis or testing tools to detect and repair bugs, such as rule-based analysis and symbolic execution [2]. However, these methods require extensive manual verification due to their high false-positive rates.
@@ I Introduction
Deep learning-based methods for code vulnerability detection are primarily divided into two types: sequence-based and graph-based approaches. Sequence-based approaches process the source code or its structures (e.g., Abstract Syntax Trees, AST) into serialized forms and interpret individual elements as tokens, which could be entire lines of code or segments divided by spaces [4, 5]. Neural networks like RNNs, LSTMs [6, 7], GRUs, and CNNs [8] are employed for detecting and classifying vulnerabilities by extracting sequence features from the code. Although sequence-based approaches exhibit strengths in learning the contextual information of code, they fall short in effectively capturing the program’s hierarchical structures, execution flows, and data and control dependencies.
Graph-based methods transform source code into heterogeneous graph structures, such as AST, Control Flow Graph (CFG), and Program Dependence Graph (PDG), to efficiently capture both local structures and dependencies within the code. These graphical representations enrich the analysis by providing intricate syntactic and semantic connections beyond mere code sequences. Leveraging code graphs, models based on GNN have demonstrated their effectiveness in extracting structural insights for vulnerability detection, as evidenced by research conducted by Wang et al. [9] and Zhou et al. [6]. Although graph-based methods provide valuable insights, they often overlook subtle coding patterns and long-distance contexts, and with their process of abstraction potentially leading to the loss of specific logic and behaviors in the code.
To address current challenges, we propose Vul-LMGNN, a novel vulnerability detection approach that combines the strengths of both pre-trained code language models (codePLM) and GNN. Vul-LMGNN constructs a code property graph (CPG) that merges ASTs, CFGs, and Program Dependency Graphs, initializing node embeddings with a pre-trained codeBERT, and utilizes a Gated Gated Neural Network (GGNN) for vulnerability detection. By jointly training codeBERT with GGNN, the proposed method implicitly fuses contextual information from code sequences with diverse information within the code property graph. Our contributions in this paper are as follows:
@@
    The proposed approach achieves state-of-the-art performance across four public datasets, outperforming previous methods. Notably, it achieves an ~10% higher F1 score on small-scale datasets.
@@
    We introduce the Gated Code GNN, which leverages a gating mechanism to capture dependency information within the code property graph, thereby effectively aggregating syntax, control flow, and data flow information.
@@
    We propose a joint training method that combines pre-trained code models with Gated GNNs, successfully capturing the benefits of both code sequence and property graph.
@@
    We introduce an auxiliary classifier designed to enhance our proposed Vul-LMGNN model by integrating predictions using linear interpolation. This augmentation further improves Vul-LMGNN’s performance with explicit fusion of predictions from two classifiers.
The rest of this paper is structured as follows: Section II provides an overview of the background and related work. Section III outlines the composition of the dataset. Section IV delves into the design specifics of our model. Section V presents the experimental outcomes and evaluates our model’s performance relative to the baseline method across datasets. Finally, Section VI summarizes the paper and outlines directions for future research.
II Related Work
@@
II-A Sequence Based Models
@@
Current studies based on deep sequence models generally follow the process of preprocessing, vectorization, and neural network modeling [4]. In data preprocessing, the raw source code is subjected to slicing and normalization techniques, after which it is parsed into a sequence of tokens. Subsequently, these tokens are transformed into vectors suitable for neural network processing. RNN and transformer-based models are used to learn contextual information within token sequences and to make the final defect prediction. RNN-based works, such as VulDeePecker [7] and SySeVR [10], have introduced lexical analysis, which converts the source code into a more fine-grained code snippet. A potential concern with code slicing is that the extracted code representations may not encompass all vulnerable code snippets. On the contrary, transformer-based methods utilize token vectorization techniques that extract more vulnerability-aware features. Transformer-based approaches often omit code slicing and normalization strategies, opting instead to directly tokenize the source code. Guo et al. [11] introduced CodeBERT, a cross-lingual pre-trained programming language model that incorporates edge prediction and node alignment tasks during training. Additionally, GraphcodeBERT [12] utilizes data flow in the pre-training stage. They can be applied to downstream detection tasks. Other methods have adopted different tokenization strategies from the NLP domain; for instance, CodeT5 [13] uses byte-level byte-pair-encoding (BPE) [14] to segment the code into tokens, while CoTEXT [15] opts for the Sentencepiece [16] model to extract tokens. These methods have been proven to be effective.
Figure 1: Overview of the Vul-LMGNN Vulnerability Detection Framework.
II-B Graph Based Models
@@
GNN-based methods also consist of three steps: preprocessing, vectorization, and neural network modeling [17, 18, 19]. During data preprocessing, the source code is transformed into various graph representations, such as AST, CFG, PDG, and CPG [20]. Then, the nodes and edges are converted into vectors, enabling the graph to be fed into a GNN model, which can learn structural information and make the final prediction. The CPG is a comprehensive code representation that combines the abstract syntax tree, control flow graph, and program dependency graph, encapsulating both the syntactic and structural information of the source code [4]. Methods like AI4VA [21] and those proposed by Feng et al. [22] directly use the original versions of the four basic graphs as their code representations. The Devign [23] was the first to employ a GNN for code vulnerability detection tasks, incorporating Natural Code Sequence (NCS) edges into the CPG. Chakraborty et al. [24] proposed the ReVeal algorithm, which combines gated GNNs with multilayer perceptrons; FUNDED [9] introduced an enhanced AST with eight additional edge types. Unlike the aforementioned strategies that add structural information, VulSPG [25] suggests eliminating code unrelated to vulnerabilities. It performs graph slicing on the CPG to generate the SPG.
GNNs struggle to capture the contextual relationships between distantly connected nodes, a limitation that models based on the Transformer architecture can effectively overcome. This insight led to our approach of integrating pre-trained code language models with code graph models. Our method utilizes a pre-trained code language model to initialize the embeddings of nodes in the code graph, jointly training the system to transfer knowledge from pre-trained code sequences to the code GNN, thus reaping the benefits of both worlds.
TABLE I: Summary of Datasets
Dataset #Vulnerable #Non-Vul Source CWEs
DiverseVul 18,945 330,492 Snyk, Bugzilla 150
Devign 11,888 14,149 Github N/A
VDSIC 82,411 119,1955 GitHub, Debian 4
ReVeal 1664 16,505 Chrome, Debian N/A
III Dataset Review
To evaluate our proposed code vulnerability detection method and other baseline methods, it is imperative to possess a substantial quantity of both vulnerable and non-vulnerable source code, spanning a diverse range of vulnerabilities. In this paper, we have selected four public code vulnerability datasets, which include three widely-used popular datasets and one newly released comprehensive dataset. We have summarized the distribution of positive and negative samples and sources of the datasets, as well as whether they distinguish specific types of vulnerability, as shown in Table I.
@@ III Dataset Review
IV Vul-LMGNN
In this section, we provide a detailed exposition of how Vul-LMGNN integrates pre-trained code language models with GNNs to achieve both implicit and explicit fusion of information. For a clearer understanding, our explanation is divided into several sections: code representation, creation of the code property graph, node embedding initialization, the operation of the gated code GNN (including its joint training with code language models), and interpolating predictions.
@@ 14             $(V,E)\leftarrow\text{extract\_code\_property\_graph}(C)$
17                   $C_{v},S_{v}\leftarrow\textit{CodeBERT}(v.\text{fragment}(),C)$
18                   $x_{v}=\text{concat}(T_{v},C_{v})$
@@
21            $\tilde{X}=\textit{GGNN}(x_{v},E)$
22             $x_{g}=\textit{Aggregate}(\tilde{X})$
@@ 34             $\theta\leftarrow\theta-\nabla_{\theta}(L_{\text{all}})$
Algorithm 1 Vul-LMGNN: Code Vulnerability Detection
IV-A Code Representation
@@
IV-B Code Property Graphs
@@
IV-C Initializing Node Embeddings with CodeBERT
Previous methods for generating node embeddings often involved training static word embedding models like Word2vec [29] on a dataset of code snippets to produce vectors for each code token. In contrast, our approach seeks to harness the power of large-scale pre-trained code language models, drawing on arge-scale pretraining to acquire prior knowledge for initializing code graph node embeddings. This is achieved by synergistically training the pre-trained code language model and GNNs on target datasets to jointly optimize node representations. Specifically, we use the pre-trained programming language model CodeBERT [11] for initializing graph node embeddings, as depicted in Fig. 3.
@@ IV-C Initializing Node Embeddings with CodeBERT
Specifically, we start by decomposing the function into a sequence of statement sets $C={c_{1},c_{2},c_{3},\ldots,c_{n}}$, where each $c_{i}$ is directly mapped to a node $v_{i}$ within the CPG. This mapping ensures that the complex structure of a function is represented as an interconnected graph of simpler, manageable elements. Each statement set is tokenized using CodeBERT’s pretrained Byte Pair Encoding (BPE) tokenizer [30], converting the statement into a series of tokens.
IV-D Gated Code Graph Neural Network
@@
IV-E Joint Training of CodeBERT and GGNN

### v2

@@ Abstract
Code Language Models (codeLMs) and Graph Neural Networks (GNNs) are widely used in code vulnerability detection. However, a critical yet often overlooked issue is that GNNs primarily rely on aggregating information from adjacent nodes, limiting structural information transfer to single-layer updates. In code graphs, nodes and relationships typically require cross-layer information propagation to fully capture complex program logic and potential vulnerability patterns. Furthermore, while some studies utilize codeLMs to supplement GNNs with code semantic information, existing integration methods have not fully explored the potential of their collaborative effects.
To address these challenges, we introduce Vul-LMGNNs that integrates pre-trained CodeLMs with GNNs, leveraging knowledge distillation to facilitate cross-layer propagation of both code semantic knowledge and structural information. Specifically, Vul-LMGNNs utilizes Code Property Graphs (CPGs) to incorporate code syntax, control flow, and data dependencies, while employing gated GNNs to extract structural information in the CPG. To achieve cross-layer information transmission, we implement an online knowledge distillation (KD) program that enables a single student GNN to acquire structural information extracted from a simultaneously trained counterpart through an alternating training procedure. Additionally, we leverage pre-trained CodeLMs to extract semantic features from code sequences. Finally, we propose an "implicit-explicit" joint training framework to better leverage the strengths of both CodeLMs and GNNs. In the implicit phase, we utilize CodeLMs to initialize the node embeddings of each student GNN. Through online knowledge distillation, we facilitate the propagation of both code semantics and structural information across layers. In the explicit phase, we perform linear interpolation between the CodeLM and the distilled GNN to learn a late fusion model. The proposed method, evaluated across four real-world vulnerability datasets, demonstrated superior performance compared to 17 state-of-the-art approaches. Our source code can be accessed via GitHub: https://github.com/Vul-LMGNN/vul-LMGGNN.
keywords:
[…]
1 Introduction
With the rapid expansion of the open-source community, software vulnerability detection technology has become a significant concern in the software industry and cybersecurity domain. Vulnerabilities pose a threat to the integrity and availability of software and computer systems, potentially leading to privilege escalation, leakage of sensitive data, denial of service, and various other attacks, resulting in substantial economic and societal losses [1, 2]. Traditionally, developers and security engineers rely on rule-based code analysis and symbolic execution techniques [3]. Although these methods can detect vulnerabilities effectively, they often generate high false-positive rates, requiring extensive manual verification.
@@ 1 Introduction
Deep learning for code vulnerability detection primarily falls into two categories: sequence-based and graph-based approaches. Sequence-based methods convert source code or its structures into serialized forms, treating elements as tokens [5, 6]. These approaches employ various architectures such as LSTMs [7, 8], CNNs [9], or modern language models to detect vulnerabilities by learning sequential features. Graph-based methods, conversely, transform source code into heterogeneous graph structures like Abstract Syntax Trees (AST), Control Flow Graphs (CFG), and Program Dependence Graphs (PDG). These methods then employ Graph Neural Networks (GNNs) to capture both local structures and dependencies, offering richer syntactic and structural insights compared to code sequences [10, 7].
However, current work has two main limitations: First, current GNN-based methods for vulnerability detection suffer from a common weakness: the GNNs they employ are inherently limited in their ability to propagate structural knowledge across multiple layers of the graph. This limitation stems from the fundamental architecture of typical GNNs, which primarily aggregate information within a single layer during each iteration of the message-passing process. Even in multi-layer GNNs, where multiple rounds of aggregation are performed, each layer’s updates are still based on immediate neighbors, failing to directly capture long-range dependencies or complex structural patterns that span across multiple layers of the graph. This localized information processing can be particularly problematic in the context of code analysis, where vulnerabilities often arise from intricate interactions between distant parts of the code structure, requiring a more holistic understanding of the program’s architecture and data flow. Secondly, GNNs offer distinct strengths in code analysis. Sequence-based approaches excel at capturing contextual semantics, while GNNs provide valuable structural insights. However, current research rarely leverages the potential synergy between these complementary approaches.
To address these challenges, we propose Vul-LMGNNs, a joint training framework that effectively integrates codeLMs and GNNs, while resolving information transfer limitations in code graph learning. The core of Vul-LMGNNs is a two-stage fusion scheme implementing an "implicit-explicit" integration: 1) Implicit Stage: We utilize semantic embeddings extracted by codeLMs to initialize nodes in the code GNN. This approach fuses semantic information with code structural information during graph update learning, constituting the initial information integration. 2) Explicit Stage: We employ linear interpolation to combine the predictions of both model types, achieving the final model fusion. Furthermore, by integrating online knowledge distillation into the first stage of our fusion architecture, we simultaneously facilitate the effective cross-layer information propagation of both semantic and graph structural information.
The contributions of our work can be summarized as follows:
@@
    We propose a novel approach for detecting code vulnerabilities that leverages codeLMs to learn the semantics from source code sequences and employs gated GNN to grasp syntactic, control, and data dependencies from code attribute graphs.
@@
    We propose an "implicit-explicit" dual-learning framework for the systematic integration of CodeLMs and GNNs. In the implicit stage, we perform an embedding-level fusion of the two models. In the explicit stage, we employ linear interpolation for the output consolidation at the model endpoints.
@@
    We propose the incorporation of an online knowledge distillation module into the joint training framework of CodeLMs and GNNs to address the limitations of cross-layer knowledge propagation in code graphs and to further facilitate the integration of code semantic information with code structural information.
@@
    The proposed method is extensively evaluated on four public datasets, achieving state-of-the-art performance, with approximately a 10% improvement in F1 scores on two challenging imbalanced datasets compared to the previous most competitive baseline methods.
The rest of this paper is structured as follows: Section 2 provides an overview of the background and related work. Section 4 outlines the composition of the dataset. Section 3 delves into the design specifics of our model. Section 5 presents the experimental outcomes and evaluates our model’s performance relative to the baseline method across datasets. Finally, Section 6 summarizes the paper and outlines directions for future research.
2 Related Work
@@
2.1 Sequence Based Models
@@
Current studies based on deep sequence models generally follow the process of preprocessing, vectorization, and neural network modeling [5, 4, 11, 12, 13]. In data preprocessing, the raw source code is subjected to slicing and normalization techniques, after which it is parsed into a sequence of tokens. Subsequently, these tokens are transformed into vectors suitable for neural network processing. RNN and transformer-based models are used to learn contextual information within token sequences and to make the final defect prediction. RNN-based works, such as VulDeePecker [8] and SySeVR [14], have introduced lexical analysis, which converts the source code into a more fine-grained code snippet. A potential concern with code slicing is that the extracted code representations may not encompass all vulnerable code snippets. On the contrary, transformer-based methods utilize token vectorization techniques that extract more vulnerability-aware features. Transformer-based approaches often omit code slicing and normalization strategies, opting instead to directly tokenize the source code. Chen et al.evaluated the performance of several Transformer-based methods, including CodeBERT[15] and GraphCodeBERT[16], on the DiverseVul dataset for vulnerability detection. Other methods have adopted different tokenization strategies from the NLP domain; for instance, CodeT5 [17] uses byte-level byte-pair-encoding (BPE) [18] to segment the code into tokens, while CoTEXT [19] opts for the Sentencepiece [20] model to extract tokens. These methods have been proven to be effective.
Figure 1: Overview of the Vul-LMGNNs Vulnerability Detection Framework.
2.2 Graph Based Models
@@
In the field of code vulnerability identification, GNN-based approaches leverage various code representations as input data, including AST, CFG, PDG, and CPG[21, 5], and then employ graph embedding techniques or graph representation learning methods[22], such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), knowledge Graph [23]. These techniques aim to transform the complex, high-dimensional graph structures into compact, low-dimensional vector representations while preserving essential structural and semantic information [24, 25, 26, 27, 28]. Methods like AI4VA [29] and those proposed by Feng et al.[30] directly use the original versions of the four basic graphs as their code representations. The Devign [31] was the first to employ a GNN for code vulnerability detection tasks, incorporating Natural Code Sequence (NCS) edges into the CPG. Chakraborty et al.[32] proposed the ReVeal algorithm, which combines gated GNNs with multilayer perceptrons; FUNDED [10] introduced an enhanced AST with eight additional edge types. Unlike the aforementioned strategies that add structural information, VulSPG [33] suggests eliminating code unrelated to vulnerabilities. It performs graph slicing on the CPG to generate the SPG.
Sequence-based models excel at learning code semantics but struggle to capture the non-linear, hierarchical structures in code that involve control flow, nested functions, and variable dependencies. GNNs excel at learning from graph-structured data but are often constrained by single-layer structural information transfer. Moreover, when applied to code graphs, they struggle to incorporate the contextual semantic information of the original code. This insight motivated our proposal to combine pre-trained code language models, code graph models, and online knowledge distillation.
@@
3 Vul-LMGNNs
In this section, we provide a detailed explanation of the Vul-LMGNNs process and its components. For clarity, our explanation is divided into several parts:
  1. 1.
    Code Representation: We utilize two perspectives for code representation:
    • Code graph representation
    • Code sequence representation
    We describe how to process these to produce the corresponding representations.
  2. 2.
    Code Property Graph Creation: We employ code property graphs to transform source code into an abstract graph representation. We introduce this code property graph here.
  3. 3.
    Gated GNN: For learning on the code property graph, we adopt a gated graph neural network. Its advantage lies in dynamically controlling information transmission and retention.
  4. 4.
    Online Knowledge Distillation for Graph Neural Networks: This part describes how we leverage online knowledge distillation to enable cross-layer knowledge transfer in the gated GNN, thereby improving its performance.
  5. 5.
    Code Sequence Semantic Information Extraction: Here, we use pre-trained code language models to extract contextual semantic embeddings from code sequences.
  6. 6.
    Joint Training: This section outlines how we jointly train the code language model and GNN to achieve implicit and explicit information fusion. This part will integrate the modules from sections 3, 4, and 5.
@@ 14             $(V,E)\leftarrow\text{extract\_code\_property\_graph}(C)$
17                   $C_{v},S_{v}\leftarrow\textit{CodeLM}(v.\text{fragment}(),C)$
18                   $x_{v}{\color[rgb]{0,0,0}\leftarrow}\text{concat}(T_{v},C_{v})$
@@
21            $\tilde{X}{\color[rgb]{0,0,0}\leftarrow}\textit{Online-Distilled GGNN}(x_{v},E)$
22             $x_{g}{\color[rgb]{0,0,0}\leftarrow}\textit{Aggregate}(\tilde{X})$
@@ 34             $\theta\leftarrow\theta-\nabla_{\theta}(L_{\text{all}})$
Algorithm 1 Vul-LMGNNs: Code Vulnerability Detection
3.1 Code Representation
@@
3.2 Code Property Graphs
@@
Figure 4: Schematic diagram of online knowledge distillation based on knowledge alignment.
3.3 Gated Graph Neural Network
@@
Input: The parameters $\theta_{1}$ and $\theta_{2}$ of the student $S_{1}$ and $S_{2}$; The calculated local structure by LSP module; hyper-parameter $\alpha$ and label $y$
Output: $S_{1}$ and $S_{2}$ with excellent performance.
1
2Initialization parameters $\theta_{1}$ and $\theta_{2}$;
3 while epochs $\leq$ max_epoch do
       /* Training loss of $S_{1}$/
4       Obtain the structure preserving loss $\mathcal{L}_{str}^{S_{1}}$ with Equ. (4) and Equ. (5);
5       Obtain the cross-entropy loss $\mathcal{L}_{ce}^{S_{1}}(y,p_{1})$;
6       Construct the total loss for $S_{1}$ as Equ. (7);
       /* Training loss of $S_{2}$/
7       Obtain the structure preserving loss $\mathcal{L}_{str}^{S_{2}}$ with Equ. (4) and Equ. (6);
8       Obtain the cross-entropy loss: $\mathcal{L}_{ce}^{S_{2}}(y,p_{2})$;
9       Construct the total loss for $S_{2}$ as Equ. (7);
       /* Alternating training of $S_{1}$ and $S_{2}$/
10       Update the parameter $\theta_{1}$ while keeping $\theta_{2}$ fixed;
11       Update the parameter $\theta_{2}$ while keeping $\theta_{1}$ fixed;
12
13 end while
Algorithm 2 collaborative distillation
3.4 Online Knowledge Distillation for GGNNs
Knowledge Distillation (KD) is a technique in machine learning that transfers knowledge from a complex model (teacher) to a simpler model (student). Originally proposed for model compression, KD has evolved to become a powerful method for improving model performance across various domains. The methods of KD can be broadly categorized into three main approaches: 1) Soft Targets: This involves using the probability distributions over classes produced by the teacher model to train the student model to capture inter-class relationships . 2) Feature Embeddings: This approach focuses on transferring knowledge from the intermediate layers of the teacher model. It involves matching the feature representations or embeddings between the teacher and student models. This can be done through various methods such as minimizing the L2 distance between feature maps or using specially designed transformation layers. 3) Structural Knowledge: This method aims to transfer the structural information learned by the teacher model. In the context of neural networks, this could involve mimicking the attention patterns in transformer models or preserving the relational information in graph neural networks. Techniques like Local Structure Preservation (LSP) in graph neural networks fall under this category.
Local Structure Preserving (LSP) is introduced as the first knowledge distillation module specifically designed for GNNs. Its purpose is to ensure that the local structure of the graph data learned by the student model closely resembles that of the teacher model, as referenced in [35, 36]. The method calculates the similarity $L_{i,j}$ between each node $i$ and its neighboring nodes $j$, where $(i,j)\in\epsilon$, in the feature space. This similarity is computed using one of several kernel functions and then normalized using a softmax function. The kernel functions for calculating the similarity $D(z_{i},z_{j})$ are given by the following equation:
$D(z_{i},z_{j})=\begin{cases}\|z_{i}-z_{j}\|_{2}^{2}&\text{Euclidean}\\ z_{i}\cdot z_{j}&\text{Linear}\\ (z_{i}\cdot z_{j}+c)^{d}&\text{Poly}\\ e^{-\frac{1}{2\sigma}\|z_{i}-z_{j}\|^{2}}&\text{RBF}\end{cases}$ (10)
The similarity $L_{i,j}$ is further defined using a softmax-like normalization:
$L_{i,j}=\frac{e^{D(z_{i},z_{j})}}{\sum_{j|(i,j)\in\epsilon}(e^{D(z_{i},z_{j})})},$ (11)
where $z_{i}$ and $z_{j}$ represent the node features. This equation normalizes the similarities, transforming them into a probability distribution.
The local structure $L_{i}$ of node $i$is described as the probability distribution of similarities between node $i$ and the nodes in its neighborhood.
To train the student model to emulate the local structure of the teacher model in the feature space, the Kullback-Leibler (KL) divergence is employed. The loss function for this LSP is given by:
$\mathcal{L}_{lsp}=\frac{1}{N}\sum_{i=1}^{N}D_{KL}(L_{i}^{T}||L_{i}^{S}).$ (12)
LSP enables the transfer of local structural information from the teacher model to the student model. However, it requires a pre-trained teacher model. Inspired by this and [37], we use an idea of online knowledge distillation that does not require a pre-trained teacher model; instead, it relies on multiple student models learning from each other.
The core of this idea is a cross-layer knowledge distillation strategy, where one student layer is aligned with a layer at a different depth in another student model. During each round of alternate training, structure information from each student layer is transferred to the previous layer of another student model, eventually spreading the structure information across all layers.
In our work, we use two student models, $S_{1}$ and $S_{2}$, with identical GGNN architectures and $H$ intermediate layers. They use $l^{S_{1}}_{i,j}$ and $l^{S_{2}}_{i,j}$ to represent the local structure of node $j$ in the $i$-th layer of $S_{1}$ and $S_{2}$ respectively.
For $S_{1}$, the local structure of its $i$-th layer is required to match the $(i+1)$-th layer of $S_{2}$, with the final layer $H$ matching the first layer. The structure preserving loss is calculated as the sum of KL divergence losses over all layers:
$\mathcal{L}_{str}^{S_{1}}=\sum_{i=1}^{H}L_{l_{i}}^{S_{1}},\quad\mathcal{L}_{str}^{S_{2}}=\sum_{i=1}^{H}L_{l_{i}}^{S_{2}}$ (13)
where
$\mathcal{L}_{l_{i}}^{S_{1}}=\frac{1}{N}\sum_{j=1}^{N}D_{KL}(l_{i+1,j}^{S_{2}}||l_{i,j}^{S_{1}})$ (14)
$\mathcal{L}_{l_{i}}^{S_{2}}=\frac{1}{N}\sum_{j=1}^{N}D_{KL}(l_{i+1,j}^{S_{1}}||l_{i,j}^{S_{2}})$ (15)
The total loss of $S_{1}$ and $S_{2}$ is formulated as:
$\mathcal{L}_{1}=\mathcal{L}_{ce}^{S_{1}}(y,p_{1})+\alpha L_{str}^{S_{1}},\mathcal{L}_{2}=\mathcal{L}_{ce}^{S_{2}}(y,p_{2})+\alpha L_{str}^{S_{2}}$ (16)
where $\mathcal{L}_{ce}^{S_{1}/S_{2}}$ represents the cross entropy loss function, which is computed using the predictions ($p_{1}$ or $p_{2}$) from each student model and the true label $y$. The parameter $\alpha$ serves as a weighting factor, adjusting the relative importance of different loss components. As outlined in Algorithm 2, the training process alternates between the two student models, $S_{1}$ and $S_{2}$, allowing them to learn from each other iteratively.
If the student number is greater than 2, each model will capture local structure information from all remaining models. The $i$-th layer’s structure preserving loss of the $k$-th model becomes
$L^{S_{k}}_{l_{i}}=\frac{1}{M-1}\frac{1}{N}\sum_{p=1\vee p\neq k}^{M}\sum_{j=1}^{N}D_{KL}(l^{Sp}_{i+1,j}||l^{S_{k}}_{i,j})$ (17)
where $M$ is the number of student models.
Distinct from prior endeavors, our online knowledge distillation procedure interlinks code language models with Graph Neural Networks (GNNs). Consequently, it achieves not only the transference of structural knowledge across GNN layers but also facilitates the amalgamation of semantic insights from language models within the GNN framework.
3.5 Extracting Code Sequence Semantics By CodeLMs
CodeLMs are designed to understand, generate, and manipulate programming code. Similar to natural language models[38], code language models are trained on vast corpora of source code from various programming languages and open-source repositories. They learn to recognize patterns, syntax, and semantics specific to programming languages, allowing them to perform a wide range of code-related tasks. Notable examples of code language models include: CodeBERT[15], GraphCodeBERT[16], CodeT5[17], CuBERT[39], CodeGen[40] etc. These models typically use transformer-based architectures found in large language models. However, they are specifically fine-tuned on programming tasks and incorporate domain-specific knowledge about code structure and best practices. By learning semantic code representations from massive codebases, CodeLMs have great potential for identifying potential security vulnerabilities in source code.
Now, we turn our attention to the handling of code sequences, which harbor a wealth of semantic information crucial for grasping the broader context and purpose of the code. Our approach to extracting this valuable semantic information begins with the transformation of function-level source code into code sequences. This conversion process is designed to preserve the logical flow that reflect the original order of declarations, operations, and control structures.
To process these code sequences, we fine-tune pre-trained codeLMs to capture the contextual information of various code elements, reveal implicit semantic features that may not be immediately apparent from the syntactic structure alone. We select several representative code language models. CodeBERT, with its bimodal pre-training in programming and natural language, excels in understanding the relationship between code and its natural language descriptions. GraphCodeBERT extends CodeBERT by incorporating code structure information. CodeT5, as a unified pre-trained encoder-decoder model, offers versatility in both comprehension and generation tasks, which can be particularly beneficial for more complex semantic analyses.
It’s worth noting that our choice of code language model is deliberately flexible. We view these models as adaptable components in our larger framework, focusing on their utility in extracting semantic information rather than modifying the models themselves. This approach offers significant advantages: It allows easy updates as new, more advanced models become available.
Figure 5: Detection accuracy for the top-30 high-frequency CWE vulnerability types in DiverseVul. Note: CodeBertGGNN represents a basic variant of the proposed Vul-LMGNNs, specifically Vul-LMGNNs using CodeBERT as the Language Model (LM)
3.6 Joint Training of CodeLM and GNNs
We propose an "implicit-explicit" joint training framework to synergistically optimize both CodeLM and GNNs, harnessing the strengths of both models. Our method unfolds in two distinct phases:
  1. 1.
    Implicit Stage: We leverage CodeLM to initialize the node embeddings of the code GNN and simultaneously optimize both CodeLM and GNNs during the GNN’s online knowledge distillation and label learning processes. This initialization process serves a dual purpose:
    1. (a)
      It facilitates the online knowledge distillation of the gated GNN, allowing for efficient transfer of knowledge from the language model to the graph structure.
    2. (b)
      It enhances the label learning process, providing a rich semantic foundation for the GNN to build upon.
    By using CodeLM for initialization and continuous optimization, we ensure that the GNN starts with and maintains a strong semantic understanding of the code, which is then refined through its graph-based processing.
  2. 2.
    Explicit Stage: This stage involves a more direct integration of the two models. We employ a linear interpolation technique to combine the outputs of CodeLM and the CodeLM-enhanced GNNs. This allows for a flexible, weighted combination of both models’ predictions.
The implicit phase ensures that the GNN benefits from CodeLM’s deep language understanding from the outset, while the explicit phase allows for a nuanced integration of both models’ strengths in the final output.
Next, we provide a detailed description of the two stages.
3.6.1 Implicit Stage
In our work, we utilize CodeLMs for initializing graph node embeddings. Specifically, we employ three different CodeLMs: CodeBERT[15], GraphCodeBERT, and CodeT5.For the sake of simplicity and clarity in our description, we will primarily focus on the process using CodeBERT as our exemplar CodeLM. The overall procedure, as depicted in Fig. 3, involves the following steps.
We start by decomposing the function into a sequence of statement sets $C={c_{1},c_{2},c_{3},\ldots,c_{n}}$, where each $c_{i}$ is directly mapped to a node $v_{i}$ within the CPG. This mapping ensures that the complex structure of a function is represented as an interconnected graph of simpler, manageable elements. Each statement set is tokenized using CodeBERT’s pretrained Byte Pair Encoding (BPE) tokenizer [41], converting the statement into a series of tokens.